OpenStudy (anonymous):

Solve the system of equations y=2x-3 and x^2+2x-4?

OpenStudy (anonymous):

what u want as answers? value of x and y?

OpenStudy (anonymous):

Yes, please.

OpenStudy (anonymous):

A. (1,-1) B. (-1, -5) c. (1, -1) and (-1, -5) Or D. no solution

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school, I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador

OpenStudy (anonymous):

Hi, thankyou. I just need help solving these questions. They're very confusing.

OpenStudy (raden):

the first step is take : y1 = y2 or x^2+2x-4 = 2x-3 moves 2x-3 to left side : x^2+2x-4-2x+3=0 x^2-1=0 solve for x

OpenStudy (anonymous):

Oh okay, thankyou.